OpenStudy (anonymous):

Show all steps used to solve the problem. One point for steps and one point for solution. Be sure that you check your solution for correctness. 15 = 4a - 5

OpenStudy (anonymous):

get value of a by moving numerical value -5 on left side to right during that movement -5 will become +5 do addition and divide that value by 4 to get value of a

OpenStudy (anonymous):

i need a answer now!

OpenStudy (anonymous):

15+5=4a 20=4a a=20/4 a=5